OpenStudy (sand-lock53):

Will Medal & Fan!!! Plz help! Algebra test 10 problems!

OpenStudy (studio-games):

Make a new question for every one and I will help.

OpenStudy (madarauchiha):

yee same here

OpenStudy (studio-games):

So like one here then make another then another one ETC.

OpenStudy (sand-lock53):

1. What is the equation of a line that has a y-intercept of 4.5? A. y = 4.5 B. y = −4.5 C. x= 4.5 D. y= x − 4.5 2. What is the equation of the line with m = 5 that goes through the point (2, 3)? A. y – 3 = 5(x + 2) B. y– 3 = 5(x − 2) C. y+ 3 = 5(x − 2) D. y + 3 = 5(x + 2) 3. What is the equation of the line with m = 3 and b = 8.25? A. y = −3x – 8.25 B. y = 3x − 8.25 C. y = − 3x + 8.25 D. y = 3x + 8.25 4. What is the equation of a vertical line that has an x-intercept of −7? A. x = −7 B. x = 7 C. y = x + 7 D. y = x − 7 5. What is the equation of the line with a slope of that goes through the point (4, 1) in slope-intercept form? A. B. C. D. 6. What is the equation of the line that has an x-intercept of −5? A. x = −5 B. x = 5 C. y = x −5 D. y = −x + 5 7. What is the equation of a horizontal line that has a y-intercept of 3? A. y=-1/3 B. y=1/3 C. y=3 D. y=-3

OpenStudy (madarauchiha):

1 is a im pretty sure

OpenStudy (studio-games):

1 a 2 b 3 d 4 a 5 (no choices shown) 6 a 7 c

OpenStudy (studio-games):

gtg now ttyl

OpenStudy (studio-games):

Hope that helped!